### v1

@@ 4.1 Milestones in the life of a satellite galaxy
The upper-right corner of Figure 6 shows the mean present-day stellar-to-subhalo mass ratio depending on pairs of characteristic times. [^6] This should not be confused with the mean characteristic times of satellites with a given $m_{\mathrm{sub}}/m_{\star}$. The largest gradients in mean present-day stellar-to-subhalo mass ratio are seen in panels involving time of maximal subhalo mass, which is closely related to $t_{\mathrm{cent}}$ (but the former can be expected to depend less on subhalo finder details). We might expect a similar dependence on $t_{\mathrm{sat}}$ but that is not as clear due to the prevalence of early-$t_{\mathrm{sat}}$ satellites already discussed. It is also apparent that recent infallers have higher mean subhalo-to-stellar mass ratios, something we discuss at length next.
@@
In Figure 6, we also show the marginalized distributions of satellites in the most massive ($M_{\mathrm{200m}}>10^{14}\,\mathrm{M}_{\odot}$) and the least massive ($10^{13}\,\mathrm{M}_{\odot}<M_{\mathrm{200m}}<3\times 10^{13}\,\mathrm{M}_{\odot}$) clusters in our sample. There are 12 and 142 such clusters and 1,494 and 1,059 satellites in each respective sample. As already discussed, only the time of maximal stellar mass shows a significant dependence on host cluster mass.
4.2 Preprocessing stars and dark matter
@@
Figure 7: Preprocessing stars and dark matter. Ratio of infall-to-maximum dark matter (left) and stellar (right) masses for satellite galaxies, as a function of infall time. Colours indicate bins in present-day host halo mass, $M_{\mathrm{200m}}$. Solid and dashed lines show means of galaxies that fell in as centrals (“direct infallers”) and as satellites of an infalling group (“indirect infallers”), respectively. Note the different vertical scales.
@@
Figure 8: Postprocessing stars and dark matter. Left: ratio of present-day-to-infall dark matter mass as a function of time since infall, colour-coded by total subhalo mass at infall. Right: ratio of present-day-to-infall stellar mass as a function of time since infall, colour-coded by infall gas-to-stellar mass ratio. In each case, the quantity shown in the colour bar is the one that maximizes differences in behaviour. Line styles are as in Figure 7.
@@
The concept of preprocessing is usually applied to the quenching of star formation prior to infall onto the current host. It is generally accepted that preprocessing plays a central role in shaping cluster galaxy populations, with evidence found both in observations (Haines et al., 2015; Olave-Rojas et al., 2018) and simulations (Bahé et al., 2019; Pallero et al., 2022). It is also known that the domain of influence of a dark matter halo is much larger than the virial radius (Fong & Han, 2021; Gao et al., 2023; Zhou & Han, 2023). Here we discuss whether and to what extent galaxies gain or lose mass separately in the form of stars or dark matter, independent of the mechanisms, before falling into their current host.
@@
As with star formation quenching, the mass growth of a galaxy can be significantly slowed down or even stopped well before it entered its current host—in some cases, any host at all. As summarized in Table 2, maximal subhalo mass occured on average 1 Gyr before infall; as shown in Figure 6 in some cases maximal subhalo mass can occur up to 8 Gyr before infall, and in a few rare cases it even happens before a galaxy is ever labelled a satellite of any host. The left panel of Figure 7 shows this dark matter preprocessing more directly: on average, satellites in massive ($M_{\mathrm{200m}}>10^{14}\,\mathrm{M}_{\odot}$) clusters fell into their current host with 70–80% their maximal dark matter mass, and as low as 60% at $z=0$. Even satellites currently residing in low-mass clusters ($10^{13}<M_{\mathrm{200m}}/\mathrm{M}_{\odot}<10^{14}$) fell in to the present host with roughly 90% their maximal dark matter mass. [^7] These overall numbers can be extracted visually from Figure 7 by combining the solid and dashed lines accounting for uncertainties, since both sets are mutually exclusive. The curves have been omitted from the figure to avoid overcrowding. In fact, the level of preprocessing depends strongly on whether a present-day satellite fell in as a central or as a satellite in a smaller host. We use the terms “direct infall” and “indirect infall” to refer to present-day satellites that fell into their current host as central or satellite galaxies, respectively. In the latter case, the subhalo became a sub-subhalo [^8] Sub-subhaloes are identified in hbt+ as having Depth $>1$., at least temporarily. At any given time, between 30 and 40% of infallers in EAGLE do so as satellites of an infalling group, i.e., through indirect infall. Galaxies that fell in directly (i.e., as centrals) did so with almost their maximal dark matter masses. The exception is directs infallers falling into the most massive clusters, which fell in having lost almost 10% of their dark matter mass on average while being centrals. Indeed, we may expect a small level of dark matter loss even for central galaxies, because tidal stripping can act well before a central halo reaches the virial radius of its neighbour (Behroozi et al., 2014), or due e.g. to relaxation following a major merger or tidal losses from close encounters (Lee et al., 2018). Indirect infallers, in contrast, lost as much as 40–50% of their dark matter to preprocessing, on average. These trends are also consistent with results in numerical simulations (e.g., Joshi et al., 2017). As expected, the amount of dark matter preprocessing increases with cosmic time for galaxies falling into their present hosts indirectly as satellites of another host. The ratio of dark matter at infall to maximal dark matter is constant with time for direct infallers in the most massive clusters, but it increases with cosmic time for clusters with $M_{\mathrm{200m}}\lesssim 10^{14}\,\mathrm{M}_{\odot}$, as neighboring gravitational potentials get stronger. [^9] We note that some galaxies do reach maximal subhalo mass after infall. However, this is not the dominant behaviour for direct nor indirect infallers at any given time. On average, therefore, it is appropriate to interpret the lower infall subhalo mass as preprocessing rather than as late growth.
@@
Stellar mass shows the opposite trend. As shown in Figure 6, it is quite rare for a galaxy to reach maximal stellar mass before infall; only 12% of present-day satellites had $z_{\mathrm{\star,max}}>z_{\mathrm{infall}}$. In fact, for galaxies with $t_{\mathrm{infall}}^{\mathrm{lookback}}\gtrsim 3$ Gya maximal stellar mass typically occurs 2 Gyr after infall to the present host. More recent infallers cover a wide range of $t_{\mathrm{\star,max}}$, with the latter tending to happen earlier, simply because $t_{\mathrm{infall}}$ happens late. As with dark matter mass, both the maximal stellar mass and the trend with infall time depend on both (present-day) cluster mass and on whether the galaxy fell in directly or indirectly, although the latter dependence is seen only for massive ($M_{\mathrm{200m}}>10^{14}\,\mathrm{M}_{\odot})$ clusters. In particular, galaxies that entered their current host, say, 8 Gya, did so on average with 60–75% their maximal stellar mass depending on cluster mass. (We have verified that the dependence on cluster mass is more significant than that on subhalo mass.) As already discussed, maximal stellar mass by and large happens after infall; this means satellites in lower-mass clusters grew almost twice as much stars after infall than their counterparts in higher-mass clusters. However, galaxies that fell into $M_{\mathrm{200m}}>10^{14}\,\mathrm{M}_{\odot}$ clusters indirectly gained no more (on average) than 20% additional mass afterwards. The fact that this maximal gain has remained constant over the last 7–8 Gyr (i.e., since $z\simeq 1$) for indirect infallers only is another way of seeing star formation quenching by preprocessing. In contrast, the post-infall stellar mass gain continues to increase even to the present day for indirect infallers into low-mass clusters.
@@ 4.3 Stars and dark matter after infall
Figure 9: Evolution in the subhalo mass-stellar mass plane within two randomly-selected clusters in EAGLE, one in the low-mass end and the other in the high-mass end of our sample (HostHaloId 16 and 1 at $z=0$ in the hbt+ catalog, respectively). From left to right, we show the motion in the $m_{\mathrm{sub}}-m_{\star}$ plane of 38 and 50 subhaloes within each cluster, respectively, selected randomly to span the full stellar mass range explored in this work but including the 6 most massive subhaloes (including the central subhalo, whose track ends on a white star), and only selecting satellites with $t_{\mathrm{sat}}<12$ Gya (see Section 4.1). Note that subhaloes with present-day stellar masses $m_{\star}<10^{9}\,\mathrm{M}_{\odot}$ are shown here for illustration but are not part of the sample analyzed in this work. All three panels show the same subhaloes per host. As labelled in the top left, each panel shows the timespan between the birth of a subhalo and the first time they were labelled a satellite of any host (left), since then until they fell into their present host for the first time (middle), and from first infall until the present day (right), except for the track of the central subhalo which is reproduced from birth to the present day in all panels. Each subhalo’s history is colour-coded by age of the Universe. The thin solid and dash-dotted black lines show the present-day satellite and central SHSMR, respectively, for reference. Gray dashed diagonal lines show constant subhalo-to-stellar mass ratios, increasing by factors of ten from bottom to top, as labelled.
Figure 10: Pre- and post-processing of the satellite SHSMR. From top to bottom we show the SHSMR at $z=0$, at the time of infall to the present host, and at the time each galaxy was labelled a satellite for the first time. From left to right we show increasing times since infall, as labelled. Black solid lines show the SHSMR for galaxies that fell in to their present hosts as centrals while coloured lines show galaxies that fell in as satellites of another host, colour-coded by time elapsed between first being labelled a satellite and falling into the present host, for satellites residing in $M_{\mathrm{200m}}>1e14\,\mathrm{M}_{\odot}$ only. Black dashed lines with circles show the SHSMR of central galaxies at the median infall time of each sample. Shaded regions show normal uncertainties on the mean. Note the different vertical scales in each row.
5 Evolving the satellite HSMR
@@ 5.1 Evolution of the mean relation
Figure 9 illustrates the evolution of the mass content of subhaloes in detail for two randomly-selected clusters, one on the low-mass end and another on the high-mass end of our sample, in terms of the milestones described in Section 4.1. The most striking difference between satellites in either cluster is when subhaloes suffer most of the mass loss: in the more massive cluster much of the mass loss happens prior to infall when satellites resided in another host (i.e., in the middle panel), while in the lower-mass cluster most of the mass loss suffered by subhaloes happens within their current host (i.e., in the rightmost panel). Indeed, some subhaloes in the massive host lost a significant fraction of their total mass before being satellites of any host at all, and many spent a significant amount of time as satellites before falling into their current host: they fell into the massive cluster indirectly as satellites. On the contrary, most galaxies in the low-mass cluster do not have a long trail in the middle panel, i.e., $t_{\mathrm{sat}}\approx t_{\mathrm{infall}}$. Many galaxies in the low-mass cluster also kept growing their stellar mass for a significant time after infall, while this is rarely the case in the massive cluster.
@@
Once mass loss begins it happens quickly; some of the latest-falling galaxies already lost an order of magnitude in total mass within the last 2 Gyr. This may be due to a number of effects. For one, the host cluster is continuously growing over time, which means the potential encountered by late infallers is much deeper—the gravitational effects significantly stronger—than that encountered by early infallers (halo growth can be traced approximately by the growth of the central subhalo depicted in Figure 9). However, stellar mass usually continues to grow well after total mass starts to decrease, in particular for early infallers. As discussed in Section 4.1, galaxies that fell early on may have still had a significant gas reservoir, while late infallers had already lost their gas supply while satellites of previous hosts and were not able to grow their stellar mass as much. The trends of Figure 7 are manifested in the fact that most tracks are already “moving” down in the bottom middle panel—decreasing $m_{\mathrm{sub}}$ before infall to the massive cluster—and that most tracks in the right-most panels have some rightward movement (i.e., increasing $m_{\star}$ after infall), although less so in the massive cluster (bottom panel). The same behaviour has been described in the IllustrisTNG simulation by Engler et al. (2021).
@@
The impact of the combination above in the SHSMR over time is shown in Figure 10. Anticipating the discussion, we note two differences with respect to the rest of the paper, which also apply to Figure 11. First, we only show the SHSMR of satellites in massive clusters, $M_{\mathrm{200m}}>10^{14}\,\mathrm{M}_{\odot}$, to avoid cluttering. For lower-mass hosts we see a reduction in the impact of preprocessing, which for $M_{\mathrm{200m}}<5\times 10^{13}\,\mathrm{M}_{\odot}$ is approximately half that discussed below hosts, as shown in Figure 15. Second, we include all satellites regardless of their present mass in each panel, i.e., we include satellites that have gone below $m_{\star}=10^{9}\,\mathrm{M}_{\odot}$ at $z=0$ but have large enough masses at infall or when first becoming a satellite, as appropriate for each panel. This latter change makes the discussion less sensitive to artificial disruption (e.g. Han et al., 2016; Bahé et al., 2019). [^10] We also exclude galaxies with $t_{\mathrm{sat}}>12$ Gya, which reduces the statistical power of these results but, as discussed in Section 4.1, has no impact on our conclusions. The top row shows that the present-day SHSMR depends on preprocessing time but only for galaxies infalling in the last 8 Gyr. It appears as though it takes at least 2 Gyr for preprocessing to become apparent in the SHSMR at late times ($t_{\mathrm{infall}}<4$ Gya). At any rate, the dependence of the SHSMR on time since infall is at least comparable to the dependence on processing time. This is consistent, for instance, with recent results on galaxy quenching in IllustrisTNG (Park et al., 2023) and as inferred in observations of clusters at $z\sim 1$ (Werner et al., 2022). The shorter extent of the dark purple line in the top-left panel means, roughly speaking, that there are no massive satellites ($m_{\star}>10^{10}\,\mathrm{M}_{\odot}$) which became satellites for the first time in the last 4 Gyr and then fell into their current host within 1 Gyr.
@@
The middle row shows the preprocessing of the SHSMR directly as the SHSMR at the time of infall, binned in an identical manner as discussed above. Galaxies with longer preprocessing times fall into their current hosts with significantly lower SHSMRs, and the dependence with time since infall is mostly removed, as expected. Here it is clear that 1 Gyr is enough preprocessing time to significantly shift the SHSMR of galaxies falling into massive clusters (at least until roughly 4 Gya).
@@
Finally, the bottom panels of Figure 10 show the SHSMR at the time present-day satellites first became satellites of any host, i.e., $t_{\mathrm{sat}}$. The left panel shows, surprisingly, that satellite galaxies have a lower SHSMR than central galaxies even when they first become satellites. This shows that will-be satellite galaxies are influenced by the larger-scale environment differently than galaxies that will remain centrals up to the present day. There is a hint that this effect may be stronger for galaxies with shorter preprocessing times, however this is not statistically significant. The lower SHSMR of will-be satellites is even apparent at the time of maximal subhalo mass (not shown) instead of $t_{\mathrm{sat}}$, showing this is not an artifact of the labelling of a galaxy as satellite by hbt+. In words, dark matter preprocessing is more dramatic than suggested by the discussion in Section 4.2: before removing mass it halts growth. Galaxies which will become satellites in the future have already began being preprocessed by the time they reach maximal mass. This has important implications for the estimation of mass loss observationally, which we discuss in Section 6.
@@
As mentioned, Figure 15 shows the pre- and post-processing of the SHSMR for satellites currently residing in low-mass clusters. The levels of preprocessing are smaller, as may be expected, but the trends are otherwise equivalent. It is noteworthy that the SHSMR at $t_{\mathrm{sat}}$ is also lower than that of centrals at the same times.

### v2

@@ 4.1 Milestones in the life of a satellite galaxy
The upper-right corner of Figure 6 shows the mean present-day stellar-to-subhalo mass ratio depending on pairs of characteristic times. [^7] This should not be confused with the mean characteristic times of satellites with a given $m_{\mathrm{sub}}/m_{\star}$. The largest gradients in mean present-day stellar-to-subhalo mass ratio are seen in panels involving time of maximum subhalo mass, which is closely related to $t_{\mathrm{cent}}$ (but the former can be expected to depend less on subhalo finder details). We might expect a similar dependence on $t_{\mathrm{sat}}$ but that is not as clear due to the prevalence of early-$t_{\mathrm{sat}}$ satellites already discussed. It is also apparent that recent infallers have higher mean subhalo-to-stellar mass ratios, something we discuss at length next.
@@
In Figure 6, we also show the marginalized distributions of satellites in the most massive ($M_{\mathrm{200m}}>10^{14}\,\mathrm{M}_{\odot}$) and the least massive ($10^{13}\,\mathrm{M}_{\odot}<M_{\mathrm{200m}}<3\times 10^{13}\,\mathrm{M}_{\odot}$) clusters in our sample. There are 12 and 142 such clusters and 1,494 and 1,059 satellites in each respective sample. As already discussed, only the time of maximum stellar mass shows a significant dependence on host cluster mass.
4.2 Pre-processing stars and dark matter
Refer to captionRefer to caption
@@ 4.2 Pre-processing stars and dark matter
Figure 7: Pre-processing stars and dark matter. Ratio of infall to maximum dark matter (left) and stellar (right) masses for satellite galaxies, as a function of infall time. Colours indicate bins in present-day host halo mass, $M_{\mathrm{200m}}$. The solid and dashed lines show means of galaxies that fell in as centrals (direct infallers) and as satellites of an infalling group (indirect infallers), respectively. The vertical scales are different.
Refer to captionRefer to caption
@@
Figure 8: Post-processing stars and dark matter. Left: Ratio of present-day to infall dark matter mass as a function of time since infall, colour-coded by total subhalo mass at infall. Right: Ratio of present-day to infall stellar mass as a function of time since infall, colour-coded by infall gas-to-stellar mass ratio. In each case the quantity shown in the colour bar is the one that maximizes differences in behaviour. Line styles are as in Figure 7.
@@
The concept of pre-processing is usually applied to the quenching of star formation prior to infall onto the current host. It is generally accepted that pre-processing plays a central role in shaping cluster galaxy populations, with evidence found both in observations (Haines et al., 2015; Olave-Rojas et al., 2018) and simulations (Bahé et al., 2019; Pallero et al., 2022). It is also known that the domain of influence of a dark matter halo is much larger than the virial radius, with its growth region bounded by the depletion radius (Fong & Han, 2021; LiHan21; Fong22; Gao et al., 2023; Zhou & Han, 2023). Here we discuss whether and to what extent galaxies gain or lose mass separately in the form of stars or dark matter, independent of the mechanisms, before falling into their current host.
@@
As with star formation quenching, the mass growth of a galaxy can be significantly slowed down or even stopped well before it entered its current host—in some cases, any host at all. As summarized in Table 2, maximum subhalo mass occured on average 1 Gyr before infall; as shown in Figure 6 in some cases maximum subhalo mass can occur up to 8 Gyr before infall, and in a few rare cases it even happens before a galaxy is ever labelled a satellite of any host. The left panel of Figure 7 shows this dark matter pre-processing more directly: on average, satellites in massive ($M_{\mathrm{200m}}>10^{14}\,\mathrm{M}_{\odot}$) clusters fell into their current host with 70–80% their maximum dark matter mass, and as low as 60% at $z=0$. Even satellites currently residing in low-mass clusters ($10^{13}<M_{\mathrm{200m}}/\mathrm{M}_{\odot}<10^{14}$) fell in to the present host with roughly 90% their maximum dark matter mass. [^8] These overall numbers can be extracted visually from Figure 7 by combining the solid and dashed lines accounting for uncertainties, since both sets are mutually exclusive. The curves have been omitted from the figure to avoid overcrowding. In fact, the level of pre-processing depends strongly on whether a present-day satellite fell in as a central or as a satellite in a smaller host. We use the terms ‘direct infall’ and ‘indirect infall’ to refer to present-day satellites that fell into their current host as central or satellite galaxies, respectively. In the latter case, the subhalo became a sub-subhalo, [^9] Sub-subhaloes are identified in hbt+ as having Depth $>1$. at least temporarily. At any given time, between 30 and 40% of infallers in EAGLE do so as satellites of an infalling group, i.e. through indirect infall. Galaxies that fell in directly (i.e. as centrals) did so with almost their maximum dark matter masses. The exception is directs infallers falling into the most massive clusters, which fell in having lost almost 10% of their dark matter mass on average while being centrals. We may expect a small level of dark matter loss even for central galaxies, because tidal stripping can act well before a central halo reaches the virial radius of its neighbour (Behroozi et al., 2014), or due e.g. to relaxation following a major merger or tidal losses from close encounters (Lee et al., 2018). Indirect infallers, in contrast, lost as much as 40–50% of their dark matter to pre-processing, on average. These trends are also consistent with results in $N$-body simulations (e.g. Joshi et al., 2017). As expected, the amount of dark matter pre-processing increases with cosmic time for galaxies falling into their present hosts indirectly as satellites of another host. The ratio of dark matter at infall to maximum dark matter is constant with time for direct infallers in the most massive clusters, but it increases with cosmic time for clusters with $M_{\mathrm{200m}}\lesssim 10^{14}\,\mathrm{M}_{\odot}$, as neighbouring gravitational potentials get stronger. [^10] We note that some galaxies do reach maximum subhalo mass after infall. However, this is not the dominant behaviour for direct nor indirect infallers at any given time. On average, therefore, it is appropriate to interpret the lower infall subhalo mass as pre-processing rather than as late growth.
@@
Stellar mass shows the opposite trend. As shown in Figure 6, it is quite rare for a galaxy to reach maximum stellar mass before infall; only 12% of present-day satellites had $z_{\mathrm{\star,max}}>z_{\mathrm{infall}}$. In fact, for galaxies with $t_{\mathrm{infall}}^{\mathrm{lookback}}\gtrsim 3$ Gya maximum stellar mass typically occurs 2 Gyr after infall to the present host. More recent infallers cover a wide range of $t_{\mathrm{\star,max}}$, with the latter tending to happen earlier, simply because $t_{\mathrm{infall}}$ happens late. As with dark matter mass, both the maximum stellar mass and the trend with infall time depend on both (present-day) cluster mass and on whether the galaxy fell in directly or indirectly, although the latter dependence is seen only for massive ($M_{\mathrm{200m}}>10^{14}\,\mathrm{M}_{\odot})$ clusters. In particular, galaxies that entered their current host, say, 8 Gya, did so on average with 60–75% their maximum stellar mass depending on cluster mass. (We have verified that the dependence on cluster mass is more significant than that on subhalo mass.) As already discussed, maximum stellar mass by and large happens after infall; this means satellites in lower-mass clusters grew almost twice as much stars after infall than their counterparts in higher-mass clusters. However, galaxies that fell into $M_{\mathrm{200m}}>10^{14}\,\mathrm{M}_{\odot}$ clusters indirectly gained no more (on average) than 20% additional mass afterwards. The fact that this maximum gain has remained constant over the last 7–8 Gyr (i.e. since $z\simeq 1$) for indirect infallers only is another way of seeing star formation quenching by pre-processing. In contrast, the post-infall stellar mass gain continues to increase even to the present day for indirect infallers into low-mass clusters.
@@ 4.3 Stars and dark matter after infall
Figure 9: Evolution in the subhalo mass-stellar mass plane within two randomly selected clusters in EAGLE, one in the low-mass end and the other in the high-mass end of our sample (HostHaloId 16 and 1 at $z=0$ in the hbt+ catalogue, respectively). From left to right, we show the motion in the $m_{\mathrm{sub}}-m_{\star}$ plane of 38 and 50 subhaloes within each cluster, respectively, selected randomly to span the full stellar mass range explored in this work but including the six most massive subhaloes (including the central subhalo, whose track ends on a white star), and only selecting satellites with $t_{\mathrm{sat}}<12$ Gya (see Section 4.1). Subhaloes with present-day stellar masses $m_{\star}<10^{9}\,\mathrm{M}_{\odot}$ are shown here for illustration, but are not part of the sample analysed in this work. All three panels show the same subhaloes per host. As labelled in the top left, each panel shows the time span between the birth of a subhalo and the first time they were labelled a satellite of any host (left), since then until they fell into their present host for the first time (middle), and from first infall until the present day (right), except for the track of the central subhalo which is reproduced from birth to the present day in all panels. Each subhalo’s history is colour-coded by age of the Universe. The thin solid and dash-dotted black lines show the present-day satellite and central SHSMR, respectively, for reference. The grey dashed diagonal lines show constant subhalo-to-stellar mass ratios, increasing by factors of ten from bottom to top, as labelled.
Figure 10: Pre- and post-processing of the satellite SHSMR. From top to bottom are shown the SHSMR at $z=0$, at the time of infall to the present host and at the time each galaxy was labelled a satellite for the first time. From left to right are shown increasing times since infall, as labelled. The black solid lines show the SHSMR for galaxies that fell in to their present hosts as centrals, while the coloured lines show galaxies that fell in as satellites of another host, colour-coded by time elapsed between first being labelled a satellite and falling into the present host, for satellites residing in $M_{\mathrm{200m}}>10^{14}\,\mathrm{M}_{\odot}$ only. The black dashed lines with circles show the SHSMR of central galaxies at the median infall time of each sample. The shaded regions show normal uncertainties on the mean. The vertical scale is different in each row.
5 Evolving the satellite SHSMR
@@ 5.1 Evolution of the mean relation
Figure 9 illustrates the evolution of the mass content of subhaloes in detail for two randomly selected clusters, one on the low-mass end and another on the high-mass end of our sample, in terms of the milestones described in Section 4.1. The most striking difference between satellites in either cluster is when subhaloes suffer most of the mass loss: in the more massive cluster much of the mass loss happens prior to infall when satellites resided in another host (i.e. in the middle panel), while in the lower-mass cluster most of the mass loss suffered by subhaloes happens within their current host (i.e. in the rightmost panel). Some subhaloes in the massive host lost a significant fraction of their total mass before being satellites of any host at all, and many spent a significant amount of time as satellites before falling into their current host: they fell into the massive cluster indirectly as satellites. On the contrary, most galaxies in the low-mass cluster do not have a long trail in the middle panel, i.e. $t_{\mathrm{sat}}\approx t_{\mathrm{infall}}$. Many galaxies in the low-mass cluster also kept growing their stellar mass for a significant time after infall, while this is rarely the case in the massive cluster.
@@
Once mass loss begins it happens quickly; some of the latest-falling galaxies already lost an order of magnitude in total mass within the last 2 Gyr. This may be due to a number of effects. For one, the host cluster is continuously growing over time, which means the potential encountered by late infallers is much deeper—the gravitational effects significantly stronger—than that encountered by early infallers (halo growth can be traced approximately by the growth of the central subhalo depicted in Figure 9). However, stellar mass usually continues to grow well after total mass starts to decrease, in particular for early infallers. As discussed in Section 4.1, galaxies that fell early on may have still had a significant gas reservoir, while late infallers had already lost their gas supply and were not able to grow their stellar mass as much. The trends of Figure 7 are manifested in the fact that most tracks are already “moving” down in the bottom middle panel—decreasing $m_{\mathrm{sub}}$ before infall to the massive cluster—and that most tracks in the right-most panels have some rightward movement (i.e. increasing $m_{\star}$ after infall), although less so in the massive cluster (bottom panel). The same behaviour has been described in the IllustrisTNG simulation by Engler et al. (2021).
@@
The impact of the combination above in the SHSMR over time is shown in Figure 10. Anticipating the discussion, we note two differences with respect to the rest of the paper, which also apply to Figure 11. First, we only show the SHSMR of satellites in massive clusters, $M_{\mathrm{200m}}>10^{14}\,\mathrm{M}_{\odot}$, to avoid cluttering. For lower-mass hosts we see a reduction in the impact of pre-processing, which for $M_{\mathrm{200m}}<5\times 10^{13}\,\mathrm{M}_{\odot}$ hosts is approximately half that discussed below, as shown in Figure 15. Second, we include all satellites regardless of their present mass in each panel, i.e. we include satellites that have gone below $m_{\star}=10^{9}\,\mathrm{M}_{\odot}$ at $z=0$ but have large enough masses at infall or when first becoming a satellite, as appropriate for each panel. This latter change makes the discussion less sensitive to artificial disruption (e.g. Han et al., 2016; Bahé et al., 2019). [^11] We also exclude galaxies with $t_{\mathrm{sat}}>12$ Gya, which reduces the statistical power of these results but, as discussed in Section 4.1, has no impact on our conclusions. The top row shows that the present-day SHSMR depends on pre-processing time but only for galaxies infalling in the last 8 Gyr. It appears as though it takes at least 2 Gyr for pre-processing to become apparent in the SHSMR at late times ($t_{\mathrm{infall}}<4$ Gya). At any rate, the dependence of the SHSMR on time since infall is at least comparable to the dependence on pre-processing time. This is consistent, for instance, with recent results on galaxy quenching in IllustrisTNG (Park et al., 2023) and as inferred in observations of clusters at $z\sim 1$ (Werner et al., 2022). The shorter extent of the dark purple line in the top-left panel means, roughly speaking, that there are no massive satellites ($m_{\star}>10^{10}\,\mathrm{M}_{\odot}$) which became satellites for the first time in the last 4 Gyr and then fell into their current host within 1 Gyr.
@@
The middle row shows the pre-processing of the SHSMR directly as the SHSMR at the time of infall, binned in an identical manner as discussed above. Galaxies with longer pre-processing times fall into their current hosts with significantly lower SHSMRs, and the dependence with time since infall is mostly removed, as expected. Here it is clear that 1 Gyr is enough pre-processing time to significantly shift the SHSMR of galaxies falling into massive clusters (at least until roughly 4 Gya).
@@
Finally, the bottom panels of Figure 10 show the SHSMR at the time present-day satellites first became satellites of any host, i.e. $t_{\mathrm{sat}}$. The left panel shows, surprisingly, that satellite galaxies have a lower SHSMR than central galaxies even when they first become satellites. This shows that will-be satellite galaxies are influenced by the larger-scale environment differently than galaxies that will remain centrals up to the present day. There is a hint that this effect may be stronger for galaxies with shorter pre-processing times; however, this is not statistically significant. The lower SHSMR of will-be satellites is even apparent at the time of maximum subhalo mass (not shown) instead of $t_{\mathrm{sat}}$, showing this is not an artefact of the labelling of a galaxy as satellite by hbt+. In words, dark matter pre-processing is more dramatic than suggested by the discussion in Section 4.2: before removing mass it halts growth. Galaxies that will become satellites in the future have already began being pre-processed by the time they reach maximum mass. This has important implications for the estimation of mass loss observationally, which we discuss in Section 6.
@@
As mentioned, Figure 15 shows the pre- and post-processing of the SHSMR for satellites currently residing in low-mass clusters. The levels of pre-processing are smaller, as may be expected, but the trends are otherwise equivalent. It is noteworthy that the SHSMR at $t_{\mathrm{sat}}$ is also lower than that of centrals at the same times.